Find two real numbers that have a sum of 6 and a product of 4.

Answer by ankor@dixie-net.com(22740)

Find two real numbers that have a sum of 6 and a product of 4.
:
x + y = 6
y = (6-x)
and
xy = 4
substitute (6-x) for y
x(6-x) = 4
:
6x - x^2 = 4
:
0 = x^2 - 6x + 4
Use the quadratic formula to find x: a=1; b=-6; c=4
Solutions
x =5.236 then y = .764
and
x =.764 then y = 5.236
;
;
Check
5.326 * .764 = 4.00
5.326 + .764 = 6